# Near-Optimal Differentially Private $k$-Core Decomposition

### 12 Dec 2023 v1

@@ Abstract
Recent work by Dhulipala, Liu, Raskhodnikova, Shi, Shun, and Yu [${}^{+}$22] initiated the study of the $k$-core decomposition problem under differential privacy. They show that approximate $k$-core numbers can be output while guaranteeing differential privacy, while only incurring a multiplicative error of $(2+\eta)$ (for any constant $\eta>0$) and additive error of $\text{poly}(\log(n))/\varepsilon$. In this paper, we revisit this problem. Our main result is an $\varepsilon$-edge differentially private algorithm for $k$-core decomposition which outputs the core numbers with no multiplicative error and $O(\text{log}(n)/\varepsilon)$ additive error. This improves upon previous work by a factor of 2 in the multiplicative error, while giving near-optimal additive error.
With a little additional work, this implies improved algorithms for densest subgraph and low out-degree ordering under differential privacy. For low out-degree ordering, we give an $\varepsilon$-edge differentially private algorithm which outputs an implicit orientation such that the out-degree of each vertex is at most $d+O(\log{n}/{\varepsilon})$, where $d$ is the degeneracy of the graph. This improves upon the best known guarantees for the problem by a factor of $4$ and gives near-optimal additive error. For densest subgraph, we give an $\varepsilon$-edge differentially private algorithm outputting a subset of nodes that induces a subgraph of density at least ${D^{*}}/{2}-O(\text{log}(n)/\varepsilon)$, where $D^{*}$ is the density for the optimal subgraph.
@@ 1 Introduction
The purpose of this paper is two-fold. First, we seek to fill in the gap for accurate, private and efficient algorithms for $k$-core decomposition. Second, we seek to contribute to the growing but nascent body of theoretical work on private $k$-core decomposition and related problems. The problem has such widespread applications because the $k$-core decomposition is a very efficient and accurate measure of the set of important, well-connected vertices in a graph. Such a set of vertices is colloquially known as the “core” of the graph. Formally, the $k$-core of a graph is the maximal induced subgraph of $G$ where each vertex in the induced subgraph has degree at least $k$. Since this definition captures a single core, the $k$-core decomposition problem (sometimes also called the coreness problem) is to identify the core number of every vertex $v$, which is defined to be the largest $k$ for which $v$ is in the $k$-core, but not in the $(k+1)$-core. [^1] We note that the $k$-core decomposition problem in the literature is sometimes used to refer to the problem of computing a hierarchy of $k$-cores based on graph connectivity [SP16]; note that it is not possible to emit such a hierarchy privately, so we (and prior work) focus on privately emitting core numbers for every vertex. Identifying the coreness values of every vertex is an important way of classifying the structural importance of each node in the graph.
@@
Private $k$-core decomposition was introduced in the recent work of Dhulipala, Liu, Raskhodnikova, Shi, Shun, and Yu [${}^{+}$22] with differential privacy guarantees. Differential privacy [DMNS06] is the gold standard for privacy in modern data analytics. They study the problem in the very strong local edge model of differential privacy (LEDP). In particular, the local model guarantees the privacy of each individual adjacency list such that no edge is known by any additional party beyond the two vertices that are the endpoints of the edge. In other words, there is no notion of a trusted third-party in the local model of differential privacy. Although such a model provides strong privacy guarantees, it is often a restrictive model for a few reasons. First, it is inherently a distributed model where vertices send messages over several rounds. Thus, it is important that the number of rounds of communication is small (e.g. $\text{poly}(\log n)$). Another subtlety of the local model is that often the number of queries to a particular vertex needs to be limited in order to achieve good approximation guarantees; this limits the set of algorithms we may consider for the problem. By ensuring a small number of distributed rounds (and number of queries to any particular vertex), when translating such algorithms to the central model (run on a centralized server on one machine), the running time often becomes superlinear. In particular, the $k$-core decomposition algorithm of [${}^{+}$22], when directly translated to the central model, requires $\Omega(n^{2}\log n)$ time. Second, algorithms in the local model often have worse approximation guarantees than algorithms in the central model for the same problem. For the $k$-core decomposition problem, [${}^{+}$22] produces approximate core numbers for every vertex with multiplicative approximation error $2+\eta$ and $O(\text{poly}(\log n)/\varepsilon)$ additive error, for any constant $\eta>0$, in $O(\log n)$ rounds. However, as there is no lower bound showing that a $(2+\eta)$-multiplicative approximation is necessary for privacy, it was left open whether one can improve on this multiplicative approximation factor. Additionally, the additive approximation in [${}^{+}$22] requires a fairly large poly-logarithmic factor and it was also left open whether this result could be improved.
@@
While the local model provides the strongest privacy guarantees, such strong privacy guarantees are sometimes unnecessary in cases where the graph is known by some external party anyways. (Take, for example, social network companies who already have access to the private network data.) In this paper, we revisit the differentially private $k$-core decomposition problem and provide improved utility guarantees in the central model of differential privacy. The central model assumes a trusted curator which has access to the private data and carefully processes it before releasing the privatized information. We work in the differential privacy models under edge-differential privacy. Our $k$-core decomposition algorithm in particular provides better utility guarantees than the recent work of [${}^{+}$22], providing an approximation with no multiplicative factor of the coreness values of every vertex with only an $O(\log n/\varepsilon)$ additive error with high probability. Our $k$-core decomposition algorithm is based on a novel formulation of the AboveThreshold procedure from the sparse vector technique which we call the multidimensional AboveThreshold technique (MAT). Like most existing algorithms for $k$-core, our algorithm is based on iterative peeling, i.e. iteratively removing the lowest degree vertices from the graph until all remaining vertices have large enough degree. The analysis of our algorithm leverages a new analysis of this iterative peeling process based on a multidimensional version of the sparse vector technique [${}^{+}$09].
@@ 1.1 Our Contributions
Our main result is a private algorithm for $k$-core decomposition which improves both the multiplicative and additive approximation guarantees of the algorithm in [${}^{+}$22]. We cleverly adapt a standard peeling algorithm for the $k$-core decomposition to satisfy $\varepsilon$-edge differential privacy and shows that the resulting algorithm achieves a $(1,O(\log{n}/\varepsilon))$-approximation to the core numbers. Since the maximum $k$-core number is within a factor of $2$ of the densest subgraph in any graph by the Nash-Williams theorem, the additive lower bound given by [NV21, FHS22] also translates to a $\Omega(\sqrt{\log(n)/\varepsilon})$ additive error lower bound for the $k$-core decomposition problem, so our result is nearly tight.
@@
There is an algorithm which gives a $(1,O(\frac{\log(n)}{\varepsilon}))$-approximation to the core numbers that is $\varepsilon$-edge differentially private and achieves the approximation guarantees with high probability.
We also study the closely related densest subgraph problem, which we define and state our results for next. The density of a subgraph, and the densest subgraph problem are defined as follows:
@@
We note that we cannot output the exact densest subgraph without violating privacy; here it is known that $\Omega(\sqrt{\log(n)/\varepsilon})$ additive error is necessary [NV21, FHS22] for any $\varepsilon$-DP algorithm that succeeds with high probability in $n$. We define an approximate densest subgraph as follows:
@@
There is an algorithm which gives a $(2,O(\log n/\varepsilon))$-approximation for the densest subgraph problem that is $\varepsilon$-edge differentially private and achieves the approximation guarantee with high probability.
@@
There is an $\varepsilon$-edge differentially private algorithm that gives a $(1,O(\log n/\varepsilon))$-approximate low out-degree ordering of the vertices.
Furthermore, all of our algorithms can be implemented in near-linear time if we lose a multiplicative factor of $1+\eta$ for any given constant $\eta>0$. This nearly matches the best running times for sequential non-private $k$-core decomposition algorithms, while maintaining a near-optimal utility for private algorithms.
@@ 1.2 Related Work
The three problems our paper study, $k$-core decomposition, densest subgraph and low out-degree ordering, are closely related and have been studied side by side in the literature. For instance, the classic (non-private) $2$-approximation to the densest subgraph problem by Charikar [Cha00] works by performing an iterative greedy peeling process that is equivalent to computing the $k$-core decomposition, followed by post-processing. Subsequent work improved the approximation factor by using the multiplicative weight update framework [BGM14, SV20]. There has also been a large body of work on dynamic densest subgraph, with recent breakthrough results of Sawlani and Wang [SW20] and Chekuri and Quanrud [CQ22] yielding $O(\mathsf{poly}(\log n))$ time update while maintaining a $(1+\eta)$-approximate densest subgraph. The works of [${}^{+}$22, DKLV23] rely heavily on ideas from this literature.
@@
On the private side, the first results for private densest subgraph began with the work of Nguyen and Vullikanti [NV21] and Farhadi et al. [FHO21]. [NV21] give algorithms which achieve a $O(2+\eta,\log(n)\log(1/\delta)/{\varepsilon})$-approximation under $(\varepsilon,\delta)$-differential privacy. On the other hand,  [FHO21] give algorithms which achieve a $(2+\eta,O({\log^{2.5}(n)\log(1/\sigma)}/{\varepsilon}))$-approximation with probability $1-\sigma$ again under $\varepsilon$-differential privacy. More recently, [${}^{+}$22] gave a $(1+\eta,O(\mathsf{poly}(\log n)/\varepsilon))$ approximation achieving $\varepsilon$-differential privacy by adapting the muliplicative weight update distributed algorithms of [BGM14, SV20] to the private setting; however, the private algorithm is not distributed and runs in $O((n+m)\log^{3}n)$ time. Both [NV21] and [FHO21] provide lower bounds on the additive error of any private densest subgraph algorithm; in particular, the lower bound of [FHO21] states that any $\varepsilon$-DP densest subgraph algorithm achieving high-probability accuracy guarantees must have additive error $\Omega(\sqrt{\log(n)/\varepsilon})$. Our densest subgraph result yields the same multiplicative factor as [NV21, FHO21] but improves the additive factor; on the other hand, we incur a worse multiplicative factor compared with [${}^{+}$22] but a better additive factor.
@@
For $k$-core decomposition, to the best of our knowledge [${}^{+}$22] is the first work to study this problem; they give the first private algorithm achieving a $(2+\eta,O(\mathsf{poly}(\log n)/\varepsilon)$-approximation in the local edge-differential privacy setting by adapting a recently developed parallel locally-adjustable dynamic algorithm from [${}^{+}$22] to the private setting. To the best of our knowledge, their result is also the first $\varepsilon$-differential privacy result. Since their approach is based on a parallel algorithm, improving on this result, e.g., to achieve any $(2-\eta)$-approximation seems challenging due to a known $\mathsf{P}$-completeness result for degeneracy [AM84] (which reduces to $k$-core decomposition). We take a different approach in this paper, and adapt a variant of the classical peeling algorithm [MB83], which outputs exact core numbers with high probability, to the private setting. We note that this algorithm is not parallelizable, and indeed, the depth (i.e., longest chain of sequential dependencies) of this algorithm could be $\Omega(n)$ in the worst case.
@@ 2 Differential Privacy Background
We say that two graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ are edge-neighboring if $V_{1}=V_{2}$ and $|E_{1}\Delta E_{2}|=1$ (i.e., they have the same vertex set and they differ by exactly one edge). We often use the notation $G_{1}\sim G_{2}$ to denote that $G_{1},G_{2}$ are edge-neighboring.

### 29 Feb 2024 v2

@@ Abstract
Recent work by Dhulipala et al. [${}^{+}$22] initiated the study of the $k$-core decomposition problem under differential privacy via a connection between low round/depth distributed/parallel graph algorithms and private algorithms with small error bounds. They showed that one can output differentially private approximate $k$-core numbers, while only incurring a multiplicative error of $(2+\eta)$ (for any constant $\eta>0$) and additive error of $\text{poly}(\log(n))/\varepsilon$. In this paper, we revisit this problem. Our main result is an $\varepsilon$-edge differentially private algorithm for $k$-core decomposition which outputs the core numbers with no multiplicative error and $O(\text{log}(n)/\varepsilon)$ additive error. This improves upon previous work by a factor of 2 in the multiplicative error, while giving near-optimal additive error. Our result relies on a novel generalized form of the sparse vector technique, which is especially well-suited for threshold-based graph algorithms; thus, we further strengthen the connection between distributed/parallel graph algorithms and differentially private algorithms.
[…]
@@ 1 Introduction
The purpose of this paper is two-fold. First, we seek to fill in the gap for accurate, private and efficient algorithms for $k$-core decomposition. Second, we seek to contribute to the growing but nascent body of theoretical work on formalizing the connection between low complexity parallel/distributed graph algorithms and differentially private graph algorithms. The problem we study in this paper have such widespread applications because the $k$-core decomposition is a very efficient and accurate measure of the set of important, well-connected vertices in a graph. Such a set of vertices is colloquially known as the “core” of the graph. Formally, the $k$-core of a graph is the maximal induced subgraph of $G$ where each vertex in the induced subgraph has degree at least $k$. Since this definition captures a single core, the $k$-core decomposition problem (sometimes also called the coreness problem) is to identify the core number of every vertex $v$, which is defined to be the largest $k$ for which $v$ is in the $k$-core, but not in the $(k+1)$-core. [^1] We note that the $k$-core decomposition problem in the literature is sometimes used to refer to the problem of computing a hierarchy of $k$-cores based on graph connectivity [SP16]; note that it is not possible to emit such a hierarchy privately, so we (and prior work) focus on privately emitting core numbers for every vertex. Identifying the coreness values of every vertex is an important way of classifying the structural importance of each node in the graph.
@@
Private $k$-core decomposition was introduced in the recent work of Dhulipala et al. [${}^{+}$22] with differential privacy guarantees; they give an algorithm that produces approximate core numbers for every vertex with a multiplicative approximation factor of $2+\eta$ and $O(\text{poly}(\log n)/\varepsilon)$ additive error, for any constant $\eta>0$, in $O(\log n)$ rounds. Differential privacy [DMNS06] is the gold standard for privacy in modern data analytics. They study the problem in the very strong local edge model of differential privacy (LEDP). In particular, the local model guarantees the privacy of each individual adjacency list such that no edge is known by any additional party beyond the two vertices that are the endpoints of the edge. In other words, there is no notion of a trusted third-party in the local model of differential privacy. Since this model is inherently a distributed model, it is particularly important that the number of rounds of communication is small (e.g. $\text{poly}(\log n)$). Another subtlety of the local model is that often the number of queries to a particular vertex needs to be limited in order to achieve good approximation guarantees; this limits the set of algorithms we may consider for the problem. By ensuring a small number of distributed rounds (and number of queries to any particular vertex), when translating such algorithms to the central model (run on a centralized server on one machine), the running time often becomes superlinear. In particular, the $k$-core decomposition algorithm of [${}^{+}$22], when directly translated to the central model, requires $\Omega(n^{2}\log n)$ time. However, as there is no lower bound showing that a $(2+\eta)$-multiplicative approximation is necessary for privacy, it was left open whether one can improve on this multiplicative approximation factor. Additionally, their additive approximation requires a fairly large poly-logarithmic factor and it was also left open whether this result could be improved.
@@
Although it is often the case that algorithms in the local model often have worse approximation guarantees than algorithms in the central model for the same problem, it is not the case in this paper. In this paper, we revisit the differentially private $k$-core decomposition problem and provide improved utility guarantees in both the central and local edge models of differential privacy. The central model assumes a trusted curator which has access to the private data and carefully processes it before releasing the privatized information. We work in the differential privacy models under edge-differential privacy. Our $k$-core decomposition algorithm in particular provides better utility guarantees than the recent work of [${}^{+}$22], providing an approximation with no multiplicative factor of the coreness values of every vertex with only an $O(\log n/\varepsilon)$ additive error with high probability. Our algorithm runs in linear $O(n+m)$ time but may use $\Omega(n)$ rounds in the distributed setting. However, we provide another algorithm in the distributed setting that further improves the additive error of the $O(\log^{2}n)$ round, $(2+\eta)$-approximation algorithm of [${}^{+}$22] from $O\left(\frac{\log^{3}n}{\varepsilon}\right)$ to $O\left(\frac{\log n}{\varepsilon}\right)$. All algorithms presented in this paper are based on a novel, simple generalized formulation of the AboveThreshold procedure from the sparse vector technique [${}^{+}$09] which we call the multidimensional AboveThreshold technique (MAT). Like most existing algorithms for $k$-core, our algorithm is based on iterative peeling, i.e. iteratively removing the lowest degree vertices from the graph until all remaining vertices have large enough degree. We show that our multidimensional version of the sparse vector technique can be easily adapted to such threshold-based algorithms. Since these threshold-based algorithms are used frequently in distributed/parallel algorithms, our paper further strengthens the link between parallel/distributed graph algorithms and differentially private algorithms via this general technique.
@@ 1.1 Our Contributions
Our main result is a private algorithm for $k$-core decomposition which improves both the multiplicative and additive approximation guarantees of the algorithm in [${}^{+}$22]. We adapt a standard peeling algorithm for the $k$-core decomposition using a clever generalization of the sparse vector technique to satisfy $\varepsilon$-edge differential privacy and shows that the resulting algorithm achieves a $(1,O(\log{n}/\varepsilon))$-approximation to the core numbers. Since the maximum $k$-core number is within a factor of $2$ of the densest subgraph in any graph by the Nash-Williams theorem, the additive lower bound given by [NV21, FHS22] also translates to a $\Omega(\sqrt{\log(n)/\varepsilon})$ additive error lower bound for the $k$-core decomposition problem, so our result is nearly tight.
@@
There is an algorithm which gives $(1,O(\frac{\log(n)}{\varepsilon}))$-approximate core numbers that is $\varepsilon$-(local) edge differentially private and guarantees the approximation with high probability.
We also give another $\varepsilon$-local edge differentially private algorithm that obtains the desired $\text{poly}(\log n)$ round complexity.
Theorem 1.4.
There is an algorithm which gives $(2+\eta,O(\frac{\log n}{\varepsilon}))$-approximate core numbers (for any constant $\eta>0$) that is $\varepsilon$-local edge differentially private, runs in $O(\log^{2}n)$ rounds, and guarantees the approximation with high probability.
In addition to $k$-core decomposition, we study the closely related densest subgraph problem, which we define and state our results for next. The density of a subgraph, and the densest subgraph problem are defined as follows:
@@
As stated above, we cannot output the exact densest subgraph without violating privacy. Thus, we define an approximate densest subgraph, where $d(H)$ denotes the density of $H$.
@@
There is an algorithm which gives a $(2,O(\log n/\varepsilon))$-approximation for the densest subgraph problem that is $\varepsilon$-(local) edge differentially private and achieves the approximation guarantee with high probability.
@@
There is an $\varepsilon$-(local) edge differentially private algorithm that gives a $(1,O(\log n/\varepsilon))$-approximate low out-degree ordering of the vertices.
As stated in our theorems, all of the algorithms in our paper can also be implemented in the local edge DP (LEDP) model with the same approximation and additive error guarantees as in the central model. [^2] For the algorithms that may take $\Omega(n)$ rounds in the local model, we put “(local)” in the theorems, since these algorithms are more suitable for the central model. Furthermore, all of our algorithms can be implemented (centrally) in near-linear time if we lose a multiplicative factor of $1+\eta$ for any given constant $\eta>0$. This nearly matches the best running times for sequential non-private $k$-core decomposition algorithms, while maintaining a near-optimal utility for private algorithms.
@@ 1.2 Related Work
The three problems our paper studies, $k$-core decomposition, densest subgraph and low out-degree ordering, are closely related and have been studied side by side in the literature. For instance, the classic (non-private) $2$-approximation to the densest subgraph problem by Charikar [Cha00] works by performing an iterative greedy peeling process. Subsequent work improved the approximation factor by using the multiplicative weight update framework [BGM14, SV20]. There has also been a large body of work on dynamic densest subgraph, with recent breakthrough results of [SW20, CQ22, ${}^{+}$24] yielding $O(\mathsf{poly}(\log n))$ update time while maintaining a $(1+\eta)$-approximate densest subgraph. The works of [${}^{+}$22, DKLV23] rely heavily on ideas from this literature.
@@
On the privacy side, the first results for private densest subgraph are the works of Nguyen and Vullikanti [NV21] and Farhadi et al. [FHO21]. [NV21] give algorithms which achieve a $O(2+\eta,\log(n)\log(1/\delta)/{\varepsilon})$-approximation under $(\varepsilon,\delta)$-differential privacy. On the other hand,  [FHO21] give algorithms which achieve a $(2+\eta,O({\log^{2.5}(n)\log(1/\sigma)}/{\varepsilon}))$-approximation with probability $1-\sigma$ under $\varepsilon$-differential privacy. More recently, [${}^{+}$22] gave a $(1+\eta,O(\mathsf{poly}(\log n)/\varepsilon))$ approximation achieving $\varepsilon$-differential privacy by adapting the muliplicative weight update distributed algorithms of [BGM14, SV20] to the private setting; however, the private algorithm is not distributed and runs in $O((n+m)\log^{3}n)$ time. Both [NV21] and [FHO21] provide lower bounds on the additive error of any private densest subgraph algorithm; in particular, their lower bounds state that any $\varepsilon$-DP densest subgraph algorithm achieving high-probability accuracy guarantees must have additive error $\Omega(\sqrt{\log(n)/\varepsilon})$. Our densest subgraph result yields the same multiplicative factor as [NV21, FHO21] but improves the additive factor; on the other hand, we incur a worse multiplicative factor compared with [${}^{+}$22] but a better additive factor.
@@
For $k$-core decomposition, to the best of our knowledge [${}^{+}$22] is the first work to study this problem; they give the first private algorithm achieving a $(2+\eta,O(\mathsf{poly}(\log n)/\varepsilon))$-approximation in the local edge-differential privacy setting by adapting a recently developed parallel locally-adjustable dynamic algorithm from [${}^{+}$22]. Since their approach is based on a parallel algorithm, improving on this result, e.g., to achieve any $(2-\eta)$-approximation seems challenging due to a known $\mathsf{P}$-completeness result for degeneracy [AM84] (which reduces to $k$-core decomposition). We take a different approach in this paper, and adapt a variant of the classical peeling algorithm [MB83], which outputs exact core numbers with high probability. We note that this algorithm is not parallelizable, and indeed, the depth (i.e., longest chain of sequential dependencies) of this algorithm could be $\Omega(n)$ in the worst case.
@@ 2 Differential Privacy Background
We say that two graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ are edge-neighboring if $V_{1}=V_{2}$ and $|E_{1}\oplus E_{2}|=1$ (i.e., they have the same vertex set and they differ by exactly one edge). We often use the notation $G_{1}\sim G_{2}$ to denote that $G_{1},G_{2}$ are edge-neighboring.
@@
2.1 Local Edge Differential Privacy
The local model of differential privacy studies the setting where all private information are kept private by the individual parties, i.e. there exists no trusted curator. For graphs, the local model is called local edge differential privacy as defined in [${}^{+}$22]. We now define this formally.
Definition 2.5.
An $\varepsilon$-local randomizer $R:\mathbf{a}\rightarrow\mathcal{Y}$ for node $v$ is an $\varepsilon$-edge differentially private algorithm that takes as input the set of its neighbors $N(v)$, represented by an adjacency list $\mathbf{a}=(b_{1},\dots,b_{|N(v)|})$.
In other words, a local randomizer for node $v$ is an edge-private algorithm which is run on the neighborhood of $v$. The information released via local randomizers is public to all nodes and the curator. The curator performs some computation on the released information and makes the result public. The overall computation is formalized via the notion of the transcript.
Definition 2.6.
A transcript $\pi$ is a vector consisting of 4-tuples $(S^{t}_{U},S^{t}_{R},S^{t}_{\varepsilon},S^{t}_{Y})$ – encoding the set of parties chosen, set of local randomizers assigned, set of randomizer privacy parameters, and set of randomized outputs produced – for each round $t$. Let $S_{\pi}$ be the collection of all transcripts and $S_{R}$ be the collection of all randomizers. Let STOP denote a special character indicating that the local randomizer’s computation halts. A protocol is an algorithm $\mathcal{A}:S_{\pi}\to(2^{[n]}\times 2^{S_{R}}\times 2^{\mathbb{R}^{\geq 0}}\times 2^{\mathbb{R}^{\geq 0}})\;\cup\{\text{STOP}\}$ mapping transcripts to sets of parties, randomizers, and randomizer privacy parameters. The length of the transcript, as indexed by $t$, is its round complexity. Given $\varepsilon>0$, a randomized protocol $\mathcal{A}$ on graph $G$ is $\varepsilon$-local edge differentially private if the algorithm that outputs the entire transcript generated by $\mathcal{A}$ is $\varepsilon$-edge differentially private on graph $G.$
The definition is difficult to parse, but it naturally corresponds with the intuition of dealing with an untrusted curator (see also the discussion in [DKLV23]). At the beginning, the curator only knows the node set $V$ and each node knows its neighborhood. In each round, a subset of the nodes perform some computation using local randomizers on their local edge information, the outputs from previous rounds, and the public information. The nodes then output something which can be seen by all other nodes and the curator. This released information is differentially private since it is released by local randomizers. It is the curator’s choice for whom to query, what local randomizer they use, and what privacy parameters the randomizer uses. Since the curator’s choice only depends on the transcripts from the previous rounds, this is exactly the definition of a protocol in the above definition.